OpenStudy (anonymous):

Which is the equation of a line that passes through the point (3, -4) and has a slope of 2.

OpenStudy (anonymous):

A) Y=2x-11 B)Y-2x-10 C)Y=2x-4 D)Y=2x-2

OpenStudy (anonymous):

they are probably asking you for the slope intercept equation. this is given by y = mx + b where m is the slope and b is the y-intercept. so you would plug 2 in for the slope and get y = 2x + b now you can plug in your points.

OpenStudy (anonymous):

So The Answer Is B ?

OpenStudy (anonymous):

for (3, -4) 3 is the x term and -4 is the y term. -4 = 2 x 3 + b -4 = 6 + b b = 2 so the final equation is y = 2x - 2

OpenStudy (anonymous):

No D

OpenStudy (anonymous):

Are you a freshman? Because this seems like Algebra 1.